if an M-16 is fired straight upward, then the height h(t) of the bullet in feet at time t in seconds is given by h(t) = -16t^2 + 325t
a. what is the height of the bullet 5 seconds after it is fired ? evaluate h(5)
b. How long does it take for the bullet to return to earth ?

Answer by lwsshak3(11628)

if an M-16 is fired straight upward, then the height h(t) of the bullet in feet at time t in seconds is given by h(t) = -16t^2 + 325t
a. what is the height of the bullet 5 seconds after it is fired ? evaluate h(5)
h(5) = -16*5^2 + 325*5
h(5) = -400 + 325*5
h(5) = 1225 ft (height of bullet 5 sec after it is fired)
..
b. How long does it take for the bullet to return to earth ?
When bullet returns to earth, h(5)=0
0 = -16t^2 + 325t
-16t^2 + 325t=0
-t(16t-325)=0
t=0 (reject, t>0)
or
16t-325=0
t=325/16
t≈20.31 sec (How long it takes for bullet to return to earth)